OpenStudy (anonymous):

What is a Real number between the square root of 13 and 14?

geerky42 (geerky42):

In other word, what number is in between \(\sqrt{13}\) and \(\sqrt{14}\)?

geerky42 (geerky42):

hmm... maybe like \(\sqrt{13.5}\) ?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

simple h under root 13 = 13..2 13.4 13.6 13.8 thats it may be it helpful for you

OpenStudy (anonymous):

thanks